OpenStudy (anonymous):

find the inverse of: f(x)=(3x)/(x+4)... can someone show me step by step???>.<

OpenStudy (anonymous):

sure solve \[x=\frac{3y}{y+4}\] for y. if this is confusing, first solve \[7=\frac{3y}{y+4}\] and keep track of the steps

OpenStudy (anonymous):

Inverse is where you exchange the x and the y. \[y=(3x)/(x+4)\] becomes: \[x = (3y)/(y+4)\] \[x(y+4)=3y\] \[xy+4x = 3y\] \[xy-3y = -4x\] \[y(x-3)=-4x\] \[y=(-4x)/(x-3)\]

OpenStudy (anonymous):

Note you need to get like terms together to pull out the y to solve for y.

OpenStudy (anonymous):

step 1 multiply both sides by \[y+4\] \[x(y+4)=3y\] step two multiply out using distributive law \[xy+4x=3y\] step three put the y's on one side of the equal sign \[4x=3y-xy\] step four factor out the y \[4x=(3-x)y\] step five divide to get y by itself \[\frac{4x}{3-x}=y\] step six write \[f^{-1}(x)=\frac{4x}{3-x}\]

OpenStudy (anonymous):

thanks for helping:),...